A company uses activity-based costing to determine the costs of its three products: A, B, and C. The budgeted cost and activity for each of the company's three activity cost pools are shown in the following table:

Activity cost pool budgeted cost product A product B product C
Activity 1 70,000 6,000 9,000 20,000
Activity 2 45,000 7,000 15,000 8,000
Activity 3 82,000 2,500 1,000 1,625

How much overhead will be assigned to product B using activity-based costing?
Answer #1

Calculate Overhead assigned to product B

Product B
Activity 1 70000/35000*9000 = 18000
Activity 2 45000/30000*15000 = 22500
Activity 3 82000/5125*1000 = 16000
Total 56500